# A parallel framework for graphical optimal transport

### v1

@@ 5.2 Algorithm
The problem in 43 is similar to 30, and we will extend the algorithm developed in Section 4 to handle general graphs. More precisely, following the derivations in Sec. 4, we can derive a method for approximating the solution of 43 similar to Algorithm 3. Note that the second set of constraints in 43 correspond to the constraints on $\Gamma$ in the tree-structured optimal transport problem 30. Special care needs to be applied to the first set of constraints. In particular, in order to derive an efficient algorithm we need the neighbour sets in Definition 4.4 to be nested as described in the following assumption.
Assumption 1.
For all $c\in{\mathcal{Q}}_{S}$ with at least two neighbors, there is a inclusion sequence $(c\cap c_{j_{1}})=(c\cap c_{j_{2}})\supseteq\dots\supseteq(c\cap c_{j_{\ell}})$ that includes all $\ell$ neighbours $c_{j_{1}},\dots,c_{j_{\ell}}$ of $c$. In case a constraint set ${\boldsymbol{\gamma}}\subset\Gamma$ is embedded in the clique $c$, then this can be added after the smallest set in the inclusion, i.e., $(c\cap c_{j_{\ell}})\supseteq{\boldsymbol{\gamma}}$.
Example 4.
We give an example of the nested structure in Assumption 1 in Fig. 9. Here, $c=\{x_{1},x_{3},x_{4}\},c_{1}=\{x_{1},x_{2},x_{3},x_{4}\},c_{2}=\{x_{1},x_{3},x_{4}\},c_{3}=\{x_{0},x_{1},x_{3}\}$, and finally ${\boldsymbol{\gamma}}=\{x_{3}\}.$
@@
1,2,3,412341,2,3,41 , 2 , 3 , 43333111144441,3,41341,3,41 , 3 , 40,1,30130,1,30 , 1 , 3
Proposition 2.
The solution to 43 is of the form ${\mathbf{B}}_{c}={\bf K}_{c}\odot{\bf U}_{c}\odot{\bf M}_{c},$ for each $c\in{\mathcal{Q}}_{C}$, where ${\bf K}_{c}=\exp(-{\bf C}_{c}/\epsilon)$ and ${\bf U}_{c}=\left({\mathbf{u}}^{c_{j}}_{i_{j}}\otimes\mathbf{1}\right)\odot\left({\mathbf{u}}^{c_{k}}_{i_{k}}\otimes\mathbf{1}\right)$ for two smaller tensors ${\mathbf{u}}^{c_{j}}_{i_{j}}$ and ${\mathbf{u}}^{c_{k}}_{i_{k}}$. These tensors are dual variables associated with the two neighbors $c_{j},c_{k}\in{\mathcal{Q}}_{S}$ of $c$, and $i_{j}$, $i_{k}$ denote the position in the corresponding inclusion sequence as defined in Assumption 1. More precisely, denoting $\ell_{c}$ the length of the inclusion sequence for each $c\in{\mathcal{Q}}_{S}$, the dual problem for (43) reads
$\max_{\begin{subarray}{c}{\mathbf{u}}^{c}_{i},i=1,\ldots,\ell_{c}\text{ for }c\in{\mathcal{Q}}_{S}\\ {\mathbf{u}}_{{\boldsymbol{\gamma}}}^{c}\text{ for }c_{\boldsymbol{\gamma}}\in{\mathcal{Q}}_{\Gamma}\\ \rho_{c}\text{ for }c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}\end{subarray}}$ $-\sum_{c\in{\mathcal{Q}}_{C}}\|{\bf K}_{c}\odot{\bf U}_{c}\odot{\bf M}_{c}\|_{1}+\sum_{c_{\boldsymbol{\gamma}}\in{\mathcal{Q}}_{\Gamma}}\langle\boldsymbol{\mu}_{\boldsymbol{\gamma}},\log({\mathbf{u}}^{c}_{{\boldsymbol{\gamma}}})\rangle+\sum_{c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}}\rho_{c}$ (44)
subject to $\sum_{i=1}^{\ell_{c}}\log({\mathbf{u}}^{c}_{i}\otimes\mathbf{1})=\log({\mathbf{u}}^{c}_{{\boldsymbol{\gamma}}}\otimes\mathbf{1}),\quad\text{for}c\in{\mathcal{Q}}_{\Gamma}$
@@
The proof of proposition 2 follows the standard Lagrangue dual theory. Recall that any separator set $c\in{\mathcal{Q}}_{S}$ has a set of neighbours $c_{1},\dots,c_{\ell}\in{\mathcal{Q}}_{C}$ that define an inclusion sequence $(c\cap c_{1})=(c\cap c_{2})\supseteq\dots\supseteq(c\cap c_{\ell})$ as in Assumption 1. To find the optimal dual variables of the dual problem (44) we propose to iterate through all $c\in{\mathcal{Q}}_{S}$, and update the dual variables ${\mathbf{u}}^{c}_{1},\dots\mathbf{u}^{c}_{\ell_{c}}$.
Due to the fourth condition in Definition 4, the modified junction tree forms a bipartite graph, when we consider the separators in ${\mathcal{Q}}_{S}$ as nodes, and the cliques in ${\mathcal{Q}}_{C}$ as edges. For example, in Fig. 7, the square node $\{0,J+1\}$ alone is one set in a bipartite graph, and the bottom gray nodes belong to another set. The bipartite structure allows for a parallelized algorithm. In each update we can apply coordinate ascent to all the variables associated with separators in one set together. Moreover, because those variables do not depend on each other, we can consider a subproblem of the dual, as we will describe in the following. Note that the iterative Bregman projection (Benamou et al. 2015) applies a similar idea to the Wasserstein barycenter problem. To be precise, for the updates related to each $c\in{\mathcal{Q}}_{S}$ it suffices to consider a subproblem of the dual (44) , which for $c\in{\mathcal{Q}}_{\Gamma}$ can be formulated as
$\max_{\begin{subarray}{c}{\mathbf{u}}_{i},i=1,\ldots,\ell\\ {\mathbf{v}}_{\ell+1}\end{subarray}}\sum_{i=1}^{\ell}-{\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}+\langle\boldsymbol{\mu}_{\boldsymbol{\gamma}},\log({\mathbf{v}}_{\ell+1})\rangle\qquad\mbox{subject to }\sum_{i=1}^{\ell}\log({\mathbf{u}}_{i}\otimes\mathbf{1})=\log({\mathbf{v}}_{\ell+1}\otimes\mathbf{1}).$ (45)
Here we have omitted the superscript of the dual variables. Moreover, we denote the dual variable ${\mathbf{u}}_{\gamma}={\mathbf{v}}_{\ell+1}$, the reason for which will become apparent in the algorithm. Also, we introduce the notation ${{\mathbf{k}}_{i}:=P_{c\cap c_{i}}({\bf K}_{c_{i}}\odot{\bf M}_{c_{i}}\odot(\bar{\mathbf{u}}_{i}\otimes\mathbf{1}))}$, where $\bar{\mathbf{u}}_{i}$ denotes the second dual variable defining the tensor ${\bf U}_{c_{i}}$ together with ${\mathbf{u}}_{i}$, as described in Proposition 2. Thus, we can write $\|{\bf K}_{c_{i}}\odot{\bf U}_{c_{i}}\odot{\bf M}_{c_{i}}\|_{1}={\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}$. Note that the case $c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}$ is a special case of (45), where ${\mathbf{v}}_{\ell+1}={\exp(\rho_{c})}\in\mathbb{R}$ is a scalar and $\boldsymbol{\mu}_{\gamma}=1$.
@@
    ${\mathbf{q}}_{2}=\left({\mathbf{k}}_{1}\odot{\mathbf{k}}_{2}\right)^{1/2};\qquad{\mathbf{q}}_{i}=\left(P_{c\cap c_{i}}({\mathbf{q}}_{i-1})^{i-1}\odot{\mathbf{k}}_{i}\right)^{1/i},\qquad i=3,\dots,\ell$ (46)
@@
    ${\mathbf{a}}_{i}=({\mathbf{v}}_{i+1}{\otimes\mathbf{1}})^{1/i}\odot{\mathbf{q}}_{i},\qquad\text{~{}~{}for }i=2,\ldots,\ell$ (47)
    ${\mathbf{v}}_{\ell+1}=(\boldsymbol{\mu}_{\gamma}./P_{\gamma}({\mathbf{q}}_{\ell}))^{\ell},\qquad{\mathbf{v}}_{i}=({\mathbf{a}}_{i}./P_{c\cap c_{i}}({\mathbf{q}}_{i-1}))^{i-1},\quad\text{for }i=3,\ldots,\ell$
@@
    ${\mathbf{u}}_{1}={\mathbf{a}}_{2}./{\mathbf{k}}_{1},\quad{\mathbf{u}}_{i}={\mathbf{a}}_{i}./{\mathbf{k}}_{i}~{}~{}\mbox{ for }i={2,\dots,\ell},$ (48)
Proposition 3 only discusses the case of $\ell>1$. In the case $\ell=1$, the constraint in (45) asserts that ${\mathbf{u}}_{1}={\mathbf{v}}_{2}$. Thus, the objective in (45) is optimized by ${\mathbf{u}}_{1}=\boldsymbol{\mu}_{\gamma}./{\mathbf{k}}_{1}$.
Remark 3.
The method can be generalized to the case where the cost cliques have a larger number of neighbors, i.e., point 3. in Definition 4 is generalized to: For all $c\in C_{C}$ it holds $|N(c)|\leq k$ for some $k\geq 2$. A parallelized algorithm for optimal transport problems on this type of junction tree can be set up by considering a $k$-partite partition, instead of the bipartite partition in the algorithm above.
@@ 6 Numerical examples
We present numerical results for two types of MOT problems. We examine the Wasserstein barycenter problem as discussed in Fan et al. (2022a, Example 1) and structured according to the graph in Fan et al. (2022a, Figure 1), and the Wasserstein Least Squares (WLS) example detailed in Sec. 2.2.2. The original graph and the modified junction tree for the WLS example are depicted in Fig. 1 and Fig. 7, respectively. For the sake of completeness, we also show modified junction tree for the Wasserstein spline example in Fig. 8. Notably, the barycenter problem is tree-structured, while the WLS problem forms a graph with a tree-width of two.
For the barycenter problem, the cost matrices $C_{(j,k)}$ in 28 are set as the squared Euclidean distances. In the WLS problem, the cost matrices follow 18 with coefficient $\alpha=10$. The constrained marginal distributions $\{\mu_{k}\}_{k\in\Gamma}$, supported on a uniform grid of $d$ points between 0 and 1, are generated from the log-normal distribution and normalized to sum to one. We set the accuracy threshold $\delta=0.2$ for the unregularized MOT problem.
Our comparison involves Algorithm 2, which solves a standard MOT problem with global regularization, and Algorithm 3, which addresses coupled bi-marginal OT with local regularization. We employ a random update rule for the global regularization method, consistent with the algorithm complexity analysis in Fan et al. (2022a). Our focus is on the number of iterations, anticipating that the parallel mechanism introduced by the local regularization method could reduce them. Each experiment is repeated five times with different random seeds, and the average number of iterations is reported in Fig. 10. Theoretical complexity bounds for the barycenter problem are also illustrated as dashed lines, with bounds for global and local regularization derived from Fan et al. (2022a, Theorem 2) and Theorem 3, respectively.
@@ 7 Conclusion
In this paper, we study the multi-marginal optimal transport (MOT) problem with graph-structured costs. We propose a novel approach to solve the MOT problem using local entropy regularization. Specifically, we reformulate the entire MOT problem as a sum of smaller MOT problems. When the underlying graph is a tree, these smaller problems involve dual variables that can be divided into two groups, forming a bipartite graph. We develop an efficient parallel algorithm that alternately updates these two groups of dual variables. A key advantage of this approach is its parallelizability, making it well-suited for GPU implementation. We also provide a rigorous complexity analysis of this approach and extend it to handle MOT problems associated with general graphs.
Appendix A Proofs of Theorems
Proof.
Proof of Theorem 1 We perform a block-coordinate descent in 32. First, minimizing 32 with respect to $u_{(j,k_{j})}$ for $j\in\Gamma$, while keeping the other variables fixed, yields the first set of updates. To minimize the objective with respect to $\rho_{j}$ and $u_{(j,k)}$ for $j\notin\Gamma$, relax the constraint with Lagrangian multiplier $\beta$, which gives the Lagrangian
$L(\{\rho_{j}\},\{u_{(j,k)}\},\{\beta_{j}\})$ $=\sum_{(j,k)\in E}\|B_{(j,k)}\|_{1}-\sum_{j\in\Gamma}\log(u_{(j,k_{j})})^{\top}\mu_{j}$ (49)
$-\sum_{j\in V\backslash\Gamma}\left(\rho_{j}+\beta_{j}^{\top}\left(\sum_{k\in N(j)}\log(u_{(j,k)})-\rho_{j}\mathbf{1}\right)\right)$ (50)

### v2

@@ 4.2 Algorithm
The problem in Eq. 30 is similar to Eq. 21, and we will extend the algorithm developed in Section 3 to handle general graphs. More precisely, following the derivations in Section 3, we can derive a method for approximating the solution of Eq. 30 similar to Algorithm 2. Note that the second set of constraints in Eq. 30 correspond to the constraints on $\Gamma$ in the tree-structured optimal transport problem Eq. 21. Special care needs to be applied to the first set of constraints. In particular, in order to derive an efficient algorithm we need the neighbour sets in Definition 4.1.4 to be nested as described in the following assumption.
{assumption}
For all $c\in{\mathcal{Q}}_{S}$ with at least two neighbors, there is a inclusion sequence $(c\cap c_{j_{1}})=(c\cap c_{j_{2}})\supseteq\dots\supseteq(c\cap c_{j_{\ell}})$ that includes all $\ell$ neighbours $c_{j_{1}},\dots,c_{j_{\ell}}$ of $c$. In case a constraint set ${\bm{\gamma}}\subset\Gamma$ is embedded in the clique $c$, then this can be added after the smallest set in the inclusion, i.e., $(c\cap c_{j_{\ell}})\supseteq{\bm{\gamma}}$.
Example 4.3.
We give an example of the nested structure in Section 4.2 in Fig. 6. Here, $c\!=\!\{x_{1},x_{3},x_{4}\}$, $c_{1}\!=\!\{x_{1},x_{2},x_{3},x_{4}\}$, $c_{2}\!=\!\{x_{1},x_{3},x_{4}\}$, $c_{3}\!=\!\{x_{0},x_{1},x_{3}\}$, and finally ${\bm{\gamma}}=\{x_{3}\}.$
@@
1,2,3,41,2,3,43311441,3,41,3,40,1,30,1,3
Proposition 4.4.
The solution to Eq. 30 is of the form ${\mathbf{B}}_{c}={\bf K}_{c}\odot{\bf U}_{c}\odot{\bf M}_{c},$ for each $c\in{\mathcal{Q}}_{C}$, where ${\bf K}_{c}=\exp(-{\bf C}_{c}/\epsilon)$ and ${\bf U}_{c}=\left({\mathbf{u}}^{c_{j}}_{i_{j}}\otimes\mathbf{1}\right)\odot\left({\mathbf{u}}^{c_{k}}_{i_{k}}\otimes\mathbf{1}\right)$ for two smaller tensors ${\mathbf{u}}^{c_{j}}_{i_{j}}$ and ${\mathbf{u}}^{c_{k}}_{i_{k}}$. These tensors are dual variables associated with the two neighbors $c_{j},c_{k}\in{\mathcal{Q}}_{S}$ of $c$, and $i_{j}$, $i_{k}$ denote the position in the corresponding inclusion sequence as defined in Assumption 4.2. More precisely, denoting $\ell_{c}$ the length of the inclusion sequence for each $c\in{\mathcal{Q}}_{S}$, the dual problem for (30) reads
(31) $\max_{\begin{subarray}{c}{\mathbf{u}}^{c}_{i},i=1,\ldots,\ell_{c}\text{ for }c\in{\mathcal{Q}}_{S}\\ {\mathbf{u}}_{{\bm{\gamma}}}^{c}\text{ for }c_{\bm{\gamma}}\in{\mathcal{Q}}_{\Gamma}\\ \rho_{c}\text{ for }c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}\end{subarray}}$ $-\sum_{c\in{\mathcal{Q}}_{C}}\|{\bf K}_{c}\odot{\bf U}_{c}\odot{\bf M}_{c}\|_{1}+\!\sum_{c_{\bm{\gamma}}\in{\mathcal{Q}}_{\Gamma}}\!\langle\bm{\mu}_{\bm{\gamma}},\log({\mathbf{u}}^{c}_{{\bm{\gamma}}})\rangle+\!\!\sum_{c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}}\!\!\rho_{c}$
subject to $\sum_{i=1}^{\ell_{c}}\log({\mathbf{u}}^{c}_{i}\otimes\mathbf{1})=\log({\mathbf{u}}^{c}_{{\bm{\gamma}}}\otimes\mathbf{1}),\quad\text{for }c\in{\mathcal{Q}}_{\Gamma}$
@@
The proof of Proposition 4.4 follows the standard Lagrangue dual theory. Recall that any separator set $c\in{\mathcal{Q}}_{S}$ has a set of neighbours $c_{1},\dots,c_{\ell}\in{\mathcal{Q}}_{C}$ that define an inclusion sequence $(c\cap c_{1})=(c\cap c_{2})\supseteq\dots\supseteq(c\cap c_{\ell})$ as in Assumption 4.2. To find the optimal dual variables of the dual problem (31) we propose to iterate through all $c\in{\mathcal{Q}}_{S}$, and update the dual variables ${\mathbf{u}}^{c}_{1},\dots\mathbf{u}^{c}_{\ell_{c}}$.
Due to the fourth condition in Definition 4.1, the modified junction tree forms a bipartite graph, when we consider the separators in ${\mathcal{Q}}_{S}$ as nodes, and the cliques in ${\mathcal{Q}}_{C}$ as edges. For example, in Fig. 4, the square node $\{0,J+1\}$ alone is one set in a bipartite graph, and the bottom gray nodes belong to another set. The bipartite structure allows for a parallelized algorithm. In each update we can apply coordinate ascent to all the variables associated with separators in one set together. Moreover, because those variables do not depend on each other, we can consider a subproblem of the dual, as we will describe in the following. Note that the iterative Bregman projection [BenCarCut15] applies a similar idea to the Wasserstein barycenter problem. To be precise, for the updates related to each $c\in{\mathcal{Q}}_{S}$ it suffices to consider a subproblem of the dual (31) , which for $c\in{\mathcal{Q}}_{\Gamma}$ can be formulated as
(32) $\max_{\begin{subarray}{c}{\mathbf{u}}_{i},i=1,\ldots,\ell\\ {\mathbf{v}}_{\ell+1}\end{subarray}}\sum_{i=1}^{\ell}-{\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}+\langle\bm{\mu}_{\bm{\gamma}},\log({\mathbf{v}}_{\ell+1})\rangle\quad\mbox{s.t. }\sum_{i=1}^{\ell}\log({\mathbf{u}}_{i}\otimes\mathbf{1})=\log({\mathbf{v}}_{\ell+1}\otimes\mathbf{1}).$
Here we have omitted the superscript of the dual variables. Moreover, we denote the dual variable ${\mathbf{u}}_{\gamma}={\mathbf{v}}_{\ell+1}$, the reason for which will become apparent in the algorithm. Also, we introduce the notation ${\mathbf{k}}_{i}:=P_{c\cap c_{i}}({\bf K}_{c_{i}}\odot{\bf M}_{c_{i}}\odot(\bar{\mathbf{u}}_{i}\otimes\mathbf{1}))$}, where $\bar{\mathbf{u}}_{i}$ denotes the second dual variable defining the tensor ${\bf U}_{c_{i}}$ together with ${\mathbf{u}}_{i}$, as described in Proposition 4.4. Thus, we can write $\|{\bf K}_{c_{i}}\odot{\bf U}_{c_{i}}\odot{\bf M}_{c_{i}}\|_{1}={\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}$. Note that the case $c\in{\mathcal{Q}}_{S}\backslash{\mathcal{Q}}_{\Gamma}$ is a special case of (32), where ${\mathbf{v}}_{\ell+1}={\exp(\rho_{c})}\in\mathbb{R}$ is a scalar and $\bm{\mu}_{\gamma}=1$.
@@
    (33) ${\mathbf{q}}_{2}=\left({\mathbf{k}}_{1}\odot{\mathbf{k}}_{2}\right)^{1/2};\qquad{\mathbf{q}}_{i}=\left(P_{c\cap c_{i}}({\mathbf{q}}_{i-1})^{i-1}\odot{\mathbf{k}}_{i}\right)^{1/i},\qquad i=3,\dots,\ell$
@@
    (34) ${\mathbf{a}}_{i}=({\mathbf{v}}_{i+1}{\otimes\mathbf{1}})^{1/i}\odot{\mathbf{q}}_{i},\qquad\text{~~for }i=2,\ldots,\ell$
    ${\mathbf{v}}_{\ell+1}=(\bm{\mu}_{\gamma}./P_{\gamma}({\mathbf{q}}_{\ell}))^{\ell},\qquad{\mathbf{v}}_{i}=({\mathbf{a}}_{i}./P_{c\cap c_{i}}({\mathbf{q}}_{i-1}))^{i-1},\quad\text{for }i=3,\ldots,\ell$
@@
    (35) ${\mathbf{u}}_{1}={\mathbf{a}}_{2}./{\mathbf{k}}_{1},\quad{\mathbf{u}}_{i}={\mathbf{a}}_{i}./{\mathbf{k}}_{i}~~\mbox{ for }i={2,\dots,\ell}.$
Proposition 4.5 only discusses the case of $\ell>1$. In the case $\ell=1$, the constraint in (32) asserts that ${\mathbf{u}}_{1}={\mathbf{v}}_{2}$. Thus, the objective in (32) is optimized by ${\mathbf{u}}_{1}=\bm{\mu}_{\gamma}./{\mathbf{k}}_{1}$.
Remark 4.6.
The method can be generalized to the case where the cost cliques have a larger number of neighbors, i.e., point 3. in Definition 4.1 is generalized to: For all $c\in C_{C}$ it holds $|N(c)|\leq k$ for some $k\geq 2$. A parallelized algorithm for optimal transport problems on this type of junction tree can be set up by considering a $k$-partite partition, instead of the bipartite partition in the algorithm above.
@@ 5 Numerical examples
We present numerical results for two types of MOT problems. We examine the Wasserstein barycenter problem as discussed in [fan2021complexity, Example 1] and structured according to the graph in [fan2021complexity, Figure 1], and the Wasserstein Least Squares (WLS) example detailed in Section 2.2.2. The original graph and the modified junction tree for the WLS example are depicted in Fig. 1 and Fig. 4, respectively. For the sake of completeness, we also show modified junction tree for the Wasserstein spline example in Fig. 5. Notably, the barycenter problem is tree-structured, while the WLS problem forms a graph with a tree-width of two.
For the barycenter problem, the cost matrices $C_{(j,k)}$ in Eq. 19 are set as the squared Euclidean distances. In the WLS problem, the cost matrices follow Eq. 16 with coefficient $\alpha=10$. The constrained marginal distributions $\{\mu_{k}\}_{k\in\Gamma}$, supported on a uniform grid of $d$ points between 0 and 1, are generated from the log-normal distribution and normalized to sum to one. We set the accuracy threshold $\delta=0.2$ for the unregularized MOT problem.
Our comparison involves Algorithm 4, which solves a standard MOT problem with global regularization, and Algorithm 2, which addresses coupled bi-marginal OT with local regularization. We employ a random update rule for the global regularization method, consistent with the algorithm complexity analysis in [fan2021complexity]. Our focus is on the number of iterations, anticipating that the parallel mechanism introduced by the local regularization method could reduce them. Each experiment is repeated five times with different random seeds, and the average number of iterations is reported in Fig. 7. Theoretical complexity bounds for the barycenter problem are also illustrated as dashed lines, with bounds for global and local regularization derived from [fan2021complexity, Theorem 2] and Theorem 3.11, respectively.
@@ 6 Conclusion
In this paper, we study the MOT problem with graph-structured costs. We propose a novel approach to solve the MOT problem using local entropy regularization. Specifically, we reformulate the entire MOT problem as a sum of smaller MOT problems. When the underlying graph is a tree, these smaller problems involve dual variables that can be divided into two groups, forming a bipartite graph. We develop an efficient parallel algorithm that alternately updates these two groups of dual variables. A key advantage of this approach is its parallelizability, making it well-suited for GPU implementation. We also provide a rigorous complexity analysis of this approach and extend it to handle MOT problems associated with general graphs.
7 Proofs of Theorems
Proof 7.1 (Proof of Theorem 3.1).
We perform a block-coordinate descent in Eq. 23. First, minimizing Eq. 23 with respect to $u_{(j,k_{j})}$ for $j\in\Gamma$, while keeping the other variables fixed, yields the first set of updates. To minimize the objective with respect to $\rho_{j}$ and $u_{(j,k)}$ for $j\notin\Gamma$, relax the constraint with Lagrangian multiplier $\beta$, which gives the Lagrangian $L(\{\rho_{j}\},\{u_{(j,k)}\},\{\beta_{j}\}=$
(36) $\sum_{(j,k)\in E}\!\!\|B_{(j,k)}\|_{1}-\sum_{j\in\Gamma}\log(u_{(j,k_{j})})^{\top}\mu_{j}-\!\!\sum_{j\in V\backslash\Gamma}\!\!\left(\rho_{j}+\beta_{j}^{\top}\!\left(\sum_{k\in N(j)}\!\!\log(u_{(j,k)})-\rho_{j}\mathbf{1}\right)\!\!\right).\!\!$